OpenStudy (godlovesme):

help ASAP!! i suck at absolute value equations .-. for which value of a does 4=a|x-4| have no solution A. -6 B. 0 C. 4 D. 6

OpenStudy (godlovesme):

@Nnesha

OpenStudy (godlovesme):

help pwease :)

Nnesha (nnesha):

sorry laptop just turned off

OpenStudy (godlovesme):

its ok c:

Nnesha (nnesha):

alright substitution would work substitute a for each option

Nnesha (nnesha):

and remember when an absolute value equal to a negative number/varaible then there would no solution

OpenStudy (godlovesme):

ok help me out with this 4= a + |-6-4| would it be 4= a -10 or 4= a+10 ?

Nnesha (nnesha):

`which value of a` so substitute `a` for each option

OpenStudy (godlovesme):

ohhhhhhhhh so i substitute for a then solve for x and plug it back in?

Nnesha (nnesha):

ye right substitute a for each option solve for x

OpenStudy (godlovesme):

ok got it thanks nnesha <3333 :*

Nnesha (nnesha):

sure ?? np \o^_^o/

Nnesha (nnesha):

what did you get ??

OpenStudy (godlovesme):

hold on

Nnesha (nnesha):

ok \o^_^o/

Nnesha (nnesha):

sorry i gtg but remember